# Deep neural networks for choice analysis: Enhancing behavioral regularity with gradient regularization

### v1

@@ Introduction
where $\mathbbm{1}\{\cdot\}$ is an indicator function that equals 1 if $\partial P_{i}(\mathbf{x}_{n})/\partial x_{d}\notin\mathbb{S}_{id}$, and 0 otherwise; and set $\mathbb{S}_{id}$ defines the expected sign of the partial derivative. Combing the mask matrix $\Psi(\mathbf{x}_{n})$ and the Jacobian matrix $\nabla\mathbf{P}(\mathbf{x}_{n})$, we define the sum-based gradient regularization using the Frobenius inner product [^1] For real matrices, we have $\langle A,B\rangle_{F}=\sum_{i,j}A_{ij}B_{ij}$.:
@@
This sum-based gradient regularization flexibly accommodates different prior assumptions on the signs of the derivatives. Set $\mathbb{S}_{id}$ can take negative values ($\mathbb{S}_{id}=\mathbb{R}^{-}$), positive values ($\mathbb{S}_{id}=\mathbb{R}^{+}$), or any real values ($\mathbb{S}_{id}=\mathbb{R}$), depending on the prior assumption on attribute $x_{d}$’s effect on demand $P_{i}(\mathbf{x}_{n})$. For example, by imposing $\mathbb{S}=\mathbb{R}^{-}$ on the direct derivatives, they are expected to be negative and penalized if non-negative. On the other hand, when there is no prior assumption regarding a derivative, we allow all possible signs by taking $\mathbb{S}=\mathbb{R}$. Despite such flexibility, we only impose negative constraints on the direct derivatives throughout our empirical experiments, which is the least controversial among all possibilities.
Alternative to the sum-based approach, we could also regularize the gradient’s magnitude, implying that demands are not expected to change drastically with small cost perturbations. Using the same notations, we define the norm-based regularization as
@@
This norm-based regularization is relatively common in the computer science literature (Drucker and Le Cun, 1991; Jakubovitz and Giryes, 2018), thus serving as a benchmark regularization method for our empirical experiments. Although smoothness is a relatively common assumption from a pure mathematical perspective, it is not founded on strong behavioral regularity beliefs due to possible threshold effects of pricing.
The regularization terms proposed above are termed as probability gradient regularizers (PGRs) because they exploit the analytical relationship between demand monotonicity and probability gradients $\nabla\mathbf{P}(\mathbf{x}_{n})$. Due to the computational chain among utilities, choice probabilities, and log-likelihoods, it is also possible to replace the probability gradients by utility and log-likelihood gradients. The two alternative regularizers are defined as:
  1. (1)
@@
In brief, by combining sum- and norm-based regularization with probability, utility, and log-likelihood gradients, we have designed six gradient regularizers. They are hereafter referred to as the sum-PGR, sum-UGR, sum-LGR, norm-PGR, norm-UGR, and norm-LGR, all of which will be tested thoroughly in our empirical experiments.
When hidden layers are not present, the norm-UGR reduces to $L_{2}$ regularization because the gradient of the utility is simply the DNN parameters $W$. On the other hand, the proposed gradient regularizers differ from commonly used sparsity regularizers (e.g., $L_{1}$ and $L_{2}$ norms) if any hidden layer is present. This is because the derivatives of our gradient regularizers become non-separable and non-linear in DNN parameters $W$, as opposed to $L_{2}$ norms whose derivatives would be separable and linear in $W$. This difference also implies that our gradient regularizers would still be effective under various averaging scheme employed in algorithms like Adam, making the proposed approaches more robust to the choice of algorithms.
@@
Our experiments use two datasets from Chicago and London, with distinguished car- and transit-dependent travel patterns, to examine the gradient regularizers and the behavioral regularity metrics. The first dataset was collected by the Chicago Metropolitan Agency for Planning in the My Daily Travel Survey in 2018–2019.222See https://www.cmap.illinois.gov/data/transportation/travel-survey. After preliminary cleaning, the full dataset retains 26,099 trips with 4 travel modes: driving, walking, train, and cycling. Around 70% of the trips use automobile, while the proportion of cycling trips is negligible. Hence the walking and cycling modes were merged into a single active mode to create a more balanced dataset. Based on the spatial information of each trip in terms of origin and destination, we compiled level of service data by utilizing Google Directions API to collect the travel time of each mode, where active times were calculated by averaging walking and cycling times. Train costs were provided by the dataset, while driving costs were computed by summing the money paid to toll plazas en route and parking lots. The $K$-nearest neighbors algorithm was applied to impute the missing data, especially for driving and train costs. This study uses in total 10 attributes: 2 continuous alternative attributes (travel time of each mode, and travel costs of driving and train), 3 discrete sociodemographics (age, household size, and number of cars in the household), and 5 sociodemographic indicators (for higher education, males, one-person households, one-car households, and high-income households, respectively). The basic statistics of the full Chicago dataset are summarized in Table A.1.
@@
The second dataset is the London Travel Demand Survey, (Hillel et al., 2018; Wang et al., 2020b) which includes 81,086 trips from 4 travel modes: walking, cycling, public transit, and driving. The proportion of cycling trips is again negligible, hence an active mode is created with travel time defined as the average between walking and cycling times. Meanwhile, the transit time is defined as the sum of access time, in-vehicle time, and transfer time. We use in total 9 attributes: 2 continuous alternative attributes (travel time of each mode, and travel costs of driving and transit), 2 discrete individual-specific variables (number of cars in the household, and number of transfers in transit), and 5 sociodemographic indicators (for the youth, the elderly, males, driving license, and one-car households, respectively). The basic statistics of the full London dataset are summarized in Table A.2.
@@
The two dataset are further reprocessed to create three sub-datasets for each to examine the effects of large versus small sample sizes, and in-domain versus out-of-domain generalizations. The first sub-dataset, named as 10K-Random, incorporates 10,000 trips with 70% randomly sampled for training, 10% for validation, and 20% for testing. This sub-dataset is considered as the benchmark for the ideal modeling scenario with sufficiently large sample size. The second sub-dataset, named as 1K-Random, includes 1,000 trips with 80% randomly sampled for training and 20% for validation. To avoid random variation due to small sample sizes, another 500 trips were randomly sampled for testing. The 1K-Random sub-dataset aims to simulate the classic choice modeling scenario where limited samples are available. By comparing the results between these two sub-datasets, we could evaluate how predictive performance and behavioral regularity vary with sample sizes. The third sub-dataset, named as 10K-Sorted, employs a different strategy for data splitting, where the 10,000 trips are sorted by driving cost, while the upper 20% were used for testing, and the lower 80% were further randomly sampled for training (70%) and validation (10%). As shown in Tables A.3 and A.4, the distributions of variables are quite different between the training and testing sets, with significantly higher mean and standard deviations in the testing set. This training-testing split scheme simulates the testing carried out on more expensive trips. Such out-of-domain generalizability is not only of theoretical interests, but also highly relevant in practice because it investigates model transferability, i.e., how the models perform in a target context distinct from their source context. One rationale of the sampling scheme for 10K-Sorted sub-datasets is that it resembles cross-city policy learning. Local governments regularly seek to implement transportation policies (e.g., congestion charging) that origin from other cities. The out-of-domain generalization can simulate the data and modeling challenges in such cross-city policy learning.
@@
Our experiments use the training set for model training, the validation set for hyperparameter searching, and the test set for model evaluation and comparison. Using the training set of 10K-Random (Chicago), we show that both Adam (Kingma and Ba, 2014) and AdamW (Loshchilov and Hutter, 2017) are empirically suitable for our experiments, while standard stochastic gradient descent (SGD) converges much slower and results in unreasonable individual demand functions (see detailed comparisons in Appendix B). The difference between Adam and AdamW lies in the implementation of weight decay as another type of regularization, which is not included in the experiments, thus teasing out the effects of gradient regularization from others. The training set was divided into 10 batches for model training, with a learning rate at $10^{-3}$. To ensure convergence, we train each model until the validation loss in consecutive iterations reaches an optimum.
Using the validation sets, we selected the optimal regularization strength $\lambda$ for each of the gradient regularizers by overall model performance, thus balancing predictive power and behavioral regularity. The DNN architecture was chosen with four hidden layers and 100 neurons per layer after random search. As the range of $\lambda$ depends on the dataset, model class, and gradient regularizer, we only show an example of the hyperparameter space in Table 1, where we took $\lambda$ values from $10^{-4}$ to $100$ in a logarithmic scale to fully demonstrate the effects of $\lambda$ and select the optimum. It is expected that the DNNs with extremely small $\lambda$’s approximate the benchmark DNN, while those with large $\lambda$’s sacrifice predictive power for behavioral regularity.
Table 1: Hyperparameter space for DNNs with sum-XGR (10K-Random, Chicago).
@@
Regularization strength $10^{-4}$, $10^{-3}$, 0.01, 0.1, 1, 10, 100
Lastly, the models are evaluated by their performance in the test sets. Particularly, to mitigate model randomness, we analyze the ensemble performance by averaging the results of 10 model replications. Section 5 will focus on comparing the model performance in the test sets, while their performance in the training and validation sets is reported in Appendix C . The models are evaluated by five metrics: log-likelihood, prediction accuracy, $F_{1}$ score, strong behavioral regularity, and weak behavioral regularity. The first three metrics focus on predictive power, measuring how well a model fits the observed outputs. Among the three metrics, log-likelihood is the most important one because of its probabilistic nature, its wide adoption in the field of discrete choice analysis, and its solid theoretical foundation for model convergence. Prediction accuracy and $F_{1}$ score are also adopted because the former is the most common metric in machine learning and the latter tackles the potential evaluation problem in imbalanced datasets. In addition to these predictive metrics, we also evaluate the models using strong and weak behavioral regularities based on Eq. 6. To empirically compute the two behavioral metrics, we set parameter $\varepsilon$ to a small negative number for strong regularity and a small positive number for weak regularity. Despite the theoretical threshold $\varepsilon=0$ for strong regularity, we set the value slightly lower than zero to enhance numerical stability and distinguish the difference between the two metrics. [^3] As a result, the empirical strong regularity of MNL presented in Section 5 could be slightly lower than 1. Section 5 will fully demonstrate the trade-off between predictive and behavioral metrics by adjusting the regularization strength.
@@
Three models are compared, including MNL models from the DCM family, standard DNNs, and a DCM-DNN hybrid model – TasteNets (Han et al., 2022). The proposed gradient regularizers are implemented on both DNNs and TasteNets. The MNL models are estimated with PyLogit, and DNNs and TasteNets are implemented with PyTorch. The following linear-in-parameter utility function is specified for the MNL models:
@@
In terms of TasteNet, we follow the model specification as in Han et al. (2022), where all taste parameters are modeled by a DNN. Specifically, our TasteNet implementation maps individual characteristics into individual-specific time and cost parameters via a feedforward NN with one hidden layer:
$V_{ni}=\tau\left(\mathbf{z}_{n};W\right)^{\top}\mathbf{x}_{ni}$ (18)
where $\tau$ represents the NN and $W$ is a set of weights. It is noteworthy that, in general, TasteNets do not guarantee behavioral regularity. In practice, however, relatively higher behavioral regularity in TasteNets could be achieved by shallower NN architecture, in conjunction with its linear and separable structure in $\mathbf{x}_{ni}$. We will show that our sum-based gradient regularizers work on both DNNs and TasteNets.
@@ Results
In this section, we present the results of our empirical work in three stages. Section 5.1 compares the behaviorally regularized DNNs and TasteNets with benchmark models, including their counterparts without regularization and MNL models, regarding predictive power and behavioral regularity metrics. Specifically, we design the large sample (10K-Random), small sample (1K-Random), and out-of-domain generalization (10K-Sorted) scenarios. We note that although sharing certain similarity, out-of-sample generalization (i.e., testing on new samples, like the 10K- and 1K-Random sub-datasets) and out-of-domain generalization (i.e., testing on new distributions, like the 10K-Sorted sub-datasets) are two different concepts. In particular, the former assumes that the training and test sets follow the same statistical pattern, whereas the latter refers to unforeseen distribution shift such as cross-city policy transfer (Liu et al., 2021). Section 5.2 further investigates how the regularization strength influences the trade-off between predictive power and behavioral regularity in each scenario. Finally, our empirical findings are summarized in Section 5.3.
@@
Using two 10K-Random sub-datasets from Chicago and London, we evaluate the regularized DNNs by five metrics in the test sets: log-likelihood, accuracy, and $F_{1}$ score that capture the models’ predictive power, as well as strong and weak regularities that describe their behavioral regularity. Table 2 summarizes the performance of six DNNs and three TasteNets with optimal regularization strengths, alongside the DNN, TasteNet, and MNL benchmarks. The model performance in the training and validation sets are also summarized in Tables C.1 and C.2, respectively. The optimal metrics across all models are marked in bold, while the model-wise optimal metrics are underlined. For DNNs and TasteNets, each metric is averaged across ten trained model replications, with standard deviations shown in parentheses. To illustrate demand monotonicity, we plot the individual demand functions of the three alternatives for selected models in Fig. 2, where light and dark curves represent the results of training replications and ensembles, respectively. Fig. 2 uses an “average individual” as the market representative, and varies the driving cost while keeping all other variables constant. By examining the large sample and in-domain scenarios (10K-Random), we have three major empirical findings.
@@
mean (std.) No GR PGR UGR LGR No GR PGR UGR LGR MNL
Log-likelihood $-1351.9$ $\mathbf{-1344.3}$ $-1350.2$ $-1347.4$ $\underline{-1438.3}$ $-1438.4$ $-1439.0$ $-1438.4$ $-1426.3$
(4.697) (5.521) (5.803) (5.110) (5.834) (6.027) (5.992) (6.099) (0)
Accuracy 0.729 0.730 0.729 0.729 0.713 0.713 0.713 0.713 0.718
(0.003) (0.002) (0.002) (0.002) (0.003) (0.002) (0.003) (0.002) (0)
$F_{1}$ score 0.691 0.698 0.694 0.696 0.654 0.654 0.654 0.654 0.669
(0.005) (0.004) (0.004) (0.004) (0.006) (0.005) (0.005) (0.005) (0)
Strong regularity 0.888 0.990 0.982 0.991 0.998 0.999 0.999 0.999 0.998
(0.066) (0.003) (0.012) (0.003) (0.002) (0.001) (0.001) (0.001) (0)
Weak regularity 0.922 0.999 0.996 0.999 0.999 1.000 1.000 1.000 1.000
(0.061) (0.001) (0.006) (0.001) (0.002) (0.001) (0.001) (0.001) (0)
Panel 2: Chicago dataset, norm-XGR
Log-likelihood $\mathbf{-1351.9}$ $-1353.9$ $-1362.0$ $-1354.0$ $\underline{-1438.3}$ $-1439.1$ $-1469.5$ $-1440.8$ $-1426.3$
(4.697) (5.018) (3.110) (4.451) (5.834) (5.836) (6.287) (5.804) (0)
Accuracy 0.729 0.729 0.725 0.727 0.713 0.713 0.710 0.712 0.718
(0.003) (0.004) (0.004) (0.003) (0.003) (0.003) (0.003) (0.003) (0)
$F_{1}$ score 0.691 0.688 0.677 0.683 0.654 0.653 0.643 0.652 0.669
(0.005) (0.007) (0.009) (0.007) (0.006) (0.005) (0.005) (0.005) (0)
Strong regularity 0.888 0.857 0.706 0.815 0.998 0.998 0.929 0.997 0.998
(0.066) (0.069) (0.051) (0.068) (0.002) (0.002) (0.03) (0.004) (0)
Weak regularity 0.922 0.893 0.756 0.851 0.999 0.999 0.941 0.998 1.000
(0.061) (0.067) (0.051) (0.069) (0.002) (0.002) (0.026) (0.003) (0)
Panel 3: London dataset, sum-XGR
Log-likelihood $-1292.0$ $\mathbf{-1288.6}$ $\mathbf{-1288.6}$ $-1305.0$ $\underline{-1305.8}$ $-1308.6$ $-1317.6$ $-1308.3$ $-1366.9$
(7.894) (9.668) (3.765) (7.316) (2.226) (2.795) (4.280) (2.670) (0)
Accuracy 0.729 0.729 0.727 0.724 0.732 0.730 0.728 0.730 0.730
(0.005) (0.004) (0.002) (0.004) (0.003) (0.002) (0.004) (0.002) (0)
$F_{1}$ score 0.727 0.728 0.725 0.721 0.728 0.726 0.723 0.726 0.726
(0.004) (0.004) (0.002) (0.006) (0.003) (0.002) (0.004) (0.002) (0)
Strong regularity 0.950 0.994 0.994 0.997 0.942 0.964 0.987 0.972 0.993
(0.034) (0.004) (0.009) (0.003) (0.021) (0.013) (0.008) (0.012) (0)
Weak regularity 0.969 0.999 0.998 1.000 0.966 1.000 1.000 1.000 1.000
(0.027) (0.001) (0.004) (0.000) (0.019) (0.000) (0.000) (0.000) (0)
@@
Figure 2: Individual demands as functions of driving cost (10K-Random, Chicago).

### v2

@@ Introduction
where $\mathbbm{1}\{\cdot\}$ is an indicator function that equals 1 if $\partial P_{i}(\mathbf{x}_{n})/\partial x_{d}\notin\mathbb{S}_{id}$, and 0 otherwise; and set $\mathbb{S}_{id}$ defines the expected sign of the partial derivative. Combining the mask matrix $\Psi(\mathbf{x}_{n})$ and the Jacobian matrix $\nabla\mathbf{P}(\mathbf{x}_{n})$, we define the sum-based gradient regularization using the Frobenius inner product: [^1] For real matrices, we have $\langle A,B\rangle_{F}=\sum_{i,j}A_{ij}B_{ij}$.
@@
This sum-based approach flexibly accommodates different prior assumptions on the signs of the derivatives. Set $\mathbb{S}_{id}$ can take negative values ($\mathbb{S}_{id}=\mathbb{R}^{-}$), positive values ($\mathbb{S}_{id}=\mathbb{R}^{+}$), or any real values ($\mathbb{S}_{id}=\mathbb{R}$), depending on the prior assumption on attribute $x_{d}$’s effect on demand $P_{i}(\mathbf{x}_{n})$. For example, by imposing $\mathbb{S}=\mathbb{R}^{-}$ on the direct derivatives, they are expected to be negative and penalized if non-negative. On the other hand, when there is no prior assumption regarding a derivative, we allow all possible signs by taking $\mathbb{S}=\mathbb{R}$. Despite such flexibility, we only impose negative constraints on the direct derivatives throughout our empirical experiments, which is the least controversial among all possibilities.
As an alternative to the sum-based approach, we could also regularize the gradient’s magnitude, implying that demands are not expected to change drastically with small cost perturbations. Using the same notations, we define the norm-based regularization as
@@
This norm-based approach is relatively common in the computer science literature (Drucker and Le Cun, 1991; Jakubovitz and Giryes, 2018), thus serving as a benchmark regularization method for our empirical experiments. Although smoothness is a relatively common assumption from a pure mathematical perspective, it is not founded on strong behavioral regularity beliefs due to possible threshold effects of pricing.
The regularization terms proposed above are termed as probability gradient regularizers (PGRs) because they exploit the analytical relationship between demand monotonicity and probability gradients $\nabla\mathbf{P}(\mathbf{x}_{n})$. Due to the computational chain among utilities, choice probabilities, and log-likelihoods, it is also possible to replace the probability gradients by utility and log-likelihood gradients. The two alternative regularizers are defined as
@@
In brief, by combining sum- and norm-based regularization with probability, utility, and log-likelihood gradients, we have designed six gradient regularizers. They are hereafter referred to as sum-PGR, sum-UGR, sum-LGR, norm-PGR, norm-UGR, and norm-LGR, all of which will be tested thoroughly in our empirical experiments.
When hidden layers are not present, norm-UGR reduces to $L_{2}$ regularization because the gradient of the utility is simply the DNN parameters $W$. On the other hand, the proposed gradient regularizers differ from commonly used sparsity regularizers (e.g., $L_{1}$ and $L_{2}$ norms) if any hidden layer is present. This is because the derivatives of our gradient regularizers become nonseparable and nonlinear in DNN parameters $W$, as opposed to $L_{2}$ norms whose derivatives would be separable and linear in $W$. This difference also implies that our gradient regularizers would still be effective under various averaging scheme employed in algorithms like Adam, making the proposed approaches more robust to the choice of algorithms.
@@
Our experiments use two datasets from Chicago and London, with distinguished car- and transit-dependent travel patterns, to examine the gradient regularizers and the behavioral regularity metrics. The first dataset was collected by the Chicago Metropolitan Agency for Planning (CMAP) in the My Daily Travel Survey in 2018–2019.222See https://www.cmap.illinois.gov/data/transportation/travel-survey. After preliminary cleaning, the full dataset retains 26,099 trips made by four travel modes: driving, walking, train (public transit), and cycling. Driving accounts for 70% of these trips, whereas the proportion of cycling trips is negligible. Hence the walking and cycling modes were merged into a single active mode to create a more balanced dataset. Based on the spatial information of each trip in terms of origin and destination, we compiled level of service data by utilizing Google Directions API to collect the travel time of each mode, where active times were calculated by averaging walking and cycling times. Train costs were provided by the dataset, while driving costs were computed by summing the money paid to toll plazas en route and parking lots. The $K$-nearest neighbors algorithm was applied to impute the missing data, especially for driving and train costs. This study uses in total ten explanatory variables: two continuous alternative attributes (time of each mode, and costs by car and train), three discrete sociodemographics (age, household size, and number of cars in the household), and five indicator variables of sociodemographics (higher education, males, one-person households, one-car households, and high-income households). The basic statistics of the full CMAP dataset are summarized in Table 5.
@@
The second dataset is the London Travel Demand Survey (LTDS; Hillel et al., 2018; Wang et al., 2020b), which has 81,086 trips made by four travel modes: walking, cycling, public transit, and driving. The proportion of cycling trips is again negligible, hence an active mode is created with travel time defined as the average of walking and cycling times. Meanwhile, the public transit time is defined as the sum of access time, in-vehicle time, and transfer time. We use in total nine explanatory variables: two continuous alternative attributes (time of each mode, and costs by car and public transit), two discrete sociodemographics (number of cars in the household, and number of public transit transfers), and five indicator variables of sociodemographics (youths, seniors, males, driving license, and one-car households). The basic statistics of the full LTDS dataset are summarized in Table 6.
@@
The two datasets were further reprocessed to create three smaller datasets for each to examine the effects of large versus small sample sizes, and in-domain versus out-of-domain generalizations. The first dataset, named as 10K-Random, incorporates 10,000 trips with 70% randomly sampled for training, 10% for validation, and 20% for testing. This dataset is considered as the benchmark for the ideal modeling scenario with a sufficiently large sample size. The second dataset, named as 1K-Random, contains 1,000 trips with 80% randomly sampled for training and 20% for validation. To avoid random variation due to small sample sizes, another 500 trips were randomly sampled for testing. The 1K-Random dataset aims to simulate a classical choice modeling scenario with only limited samples available. By comparing the results between these two datasets, we could evaluate how predictive performance and behavioral regularity vary with sample sizes. The third dataset, named as 10K-Sorted, employs a different strategy for data splitting, where the 10,000 trips are sorted by driving cost, while the upper 20% were used for testing, and the lower 80% were further randomly sampled for training (70%) and validation (10%). As shown in Section A.2, the distributions of variables are quite different between the training and test sets, with significantly higher mean and standard deviations in the test set. This train–test split scheme simulates the testing carried out on more expensive trips. Such out-of-domain generalizability is not only of theoretical interests, but also highly relevant in practice because it investigates model transferability, i.e., how the models perform in a target context distinct from their source context. One rationale of the sampling scheme for 10K-Sorted datasets is that it resembles cross-city policy learning. Local governments regularly seek to implement transportation policies (e.g., congestion charging) that originate in other cities. The out-of-domain generalization can simulate the data and modeling challenges in such cross-city policy learning.
@@
[…]
Table 1: Hyperparameter space for DNNs with sum-XGR (10K-Random, CMAP).
@@
[…]
Lastly, all models are evaluated by their performance in the test sets. Particularly, to mitigate model randomness, we analyze the ensemble performance by averaging the results of 10 model replications. Section 5 will focus on comparing model performance using the test sets, while the training and validation performance are reported in Appendix C. The models are evaluated by five metrics: log-likelihood, prediction accuracy, $F_{1}$ score, strong behavioral regularity, and weak behavioral regularity. The first three metrics focus on predictive power, measuring how well a model fits the observed outputs. Among the three metrics, log-likelihood is the most important one because of its probabilistic nature, its wide adoption in the field of discrete choice analysis, and its solid theoretical foundation for model convergence. Prediction accuracy and $F_{1}$ score are also adopted because the former is the most common metric in machine learning and the latter tackles the potential evaluation problem in imbalanced datasets. In addition to these predictive metrics, we also evaluate the models using strong and weak behavioral regularities based on Eq. 6. To empirically compute the two behavioral metrics, we set parameter $\varepsilon$ to a small negative number for strong regularity and a small positive number for weak regularity. Despite the theoretical threshold $\varepsilon=0$ for strong regularity, we set the value slightly lower than zero to enhance numerical stability and distinguish between the two metrics. [^3] As a result, the empirical strong regularity of MNL presented in Section 5 could be slightly lower than 100%. We conducted a sensitivity test in Appendix E to discuss the impact of $\varepsilon$ on the metric values. Section 5 will fully demonstrate the trade-off between predictive and behavioral metrics by adjusting the strength of regularization.
@@
Three types of models are compared, including MNL models from the DCM family, standard DNNs, and TasteNets, a hybrid model combining features of DCMs and NNs (Han et al., 2022). The proposed gradient regularizers are implemented on both DNNs and TasteNets. The MNL models are estimated with PyLogit, and DNNs and TasteNets are implemented with PyTorch. The following linear-in-parameters utility function is specified for the MNL models:
@@
For TasteNets, we follow the model specification as in Han et al. (2022). Specifically, this implementation maps individual characteristics into individual-specific taste parameters for alternative attributes via a feedforward NN with one hidden layer:
$V_{ni}=T(\mathbf{z}_{n};W)^{\top}\mathbf{x}_{ni}=\bm{\beta}_{ni}^{\top}\mathbf{x}_{ni}$ (18)
where $T$ represents the NN architecture, $W$ is a set of weights, and $\bm{\beta}_{ni}$ is a vector of taste parameters specific to individual $n$ and alternative $i$. Han et al. (2022) suggested imposing hard constraints on taste parameters to enforce behavioral regularity, given that Eq. 18 is linear in taste parameters. For example, let $b_{ni}$ denote the NN-embedded taste parameter for travel cost, we can enforce the sign constraint by setting $-\mathrm{ReLU}(-b_{ni})$ or $-\exp(-b_{ni})$ to ensure behaviorally regular TasteNets. We also compare the performance of soft and hard constraints on TasteNets in Section 5.
@@ Results
In this section, we present the results of our empirical work in three stages. Section 5.1 compares the behaviorally regularized DNNs and TasteNets with benchmark models, including their counterparts without regularization and MNL models, regarding predictive power and behavioral regularity metrics. Specifically, we design the large sample (10K-Random), small sample (1K-Random), and out-of-domain generalization (10K-Sorted) scenarios. We note that although sharing certain similarity, out-of-sample generalization (i.e., testing on new samples, like the 10K- and 1K-Random datasets) and out-of-domain generalization (i.e., testing on new distributions, like the 10K-Sorted datasets) are two different concepts. In particular, the former assumes that the training and test sets follow the same statistical pattern, whereas the latter refers to unforeseen distribution shift such as cross-city policy transfer (Liu et al., 2021). Section 5.2 further investigates how the strength of regularization influences the trade-off between predictive power and behavioral regularity in each scenario. Finally, our empirical findings are summarized in Section 5.3.
@@
Using two 10K-Random datasets from Chicago and London, we evaluate the regularized DNNs by five metrics in the test sets: log-likelihood, accuracy, and $F_{1}$ score that capture the models’ predictive power, as well as strong and weak regularities that describe their behavioral regularity. Table 2 summarizes the performance of DNNs and TasteNets with optimal regularization strengths, TasteNets with hard constraints, alongside the DNN, TasteNet, and MNL benchmarks. The training and validation performance of the models are summarized in Section C.1. In each row, the best and second best metrics are marked in bold and underlined, respectively. For DNNs and TasteNets, each metric is averaged across ten trained model replications, with standard deviations shown in parentheses. To illustrate demand monotonicity, we plot the individual demand functions of the three alternatives for selected models in Fig. 2, where light and dark curves represent the results of training replications and ensembles, respectively. Fig. 2 uses an “average individual” as the market representative, and varies the driving cost while keeping all other variables constant. Furthermore, Appendix D illustrates the impact of regularization strength on individual demands. By examining the large sample and in-domain scenarios (10K-Random), we have three major empirical findings.
@@
Mean (SD) No GR PGR UGR LGR No GR PGR UGR LGR ReLU Exp MNL
Panel 1: CMAP data, sum-XGR
Log-likelihood $-1351.9$ $\mathbf{-1344.3}$ $-1350.2$ $\underline{-1347.4}$ $-1438.3$ $-1438.4$ $-1439.0$ $-1438.4$ $-1463.0$ $-1439.7$ $-1426.3$
(4.697) (5.521) (5.803) (5.110) (5.834) (6.027) (5.992) (6.099) (20.05) (6.073) (0)
Accuracy 0.729 0.730 0.729 0.729 0.713 0.713 0.713 0.713 0.712 0.717 0.718
(0.003) (0.002) (0.002) (0.002) (0.003) (0.002) (0.003) (0.002) (0.002) (0.003) (0)
$F_{1}$ score 0.691 0.698 0.694 0.696 0.654 0.654 0.654 0.654 0.650 0.664 0.669
(0.005) (0.004) (0.004) (0.004) (0.006) (0.005) (0.005) (0.005) (0.002) (0.004) (0)
Strong regularity 0.888 0.990 0.982 0.991 0.998 0.999 0.999 0.999 0.426 0.999 0.998
(0.066) (0.003) (0.012) (0.003) (0.002) (0.001) (0.001) (0.001) (0.320) (0.001) (0)
Weak regularity 0.922 0.999 0.996 0.999 0.999 1.000 1.000 1.000 1.000 1.000 1.000
(0.061) (0.001) (0.006) (0.001) (0.002) (0.001) (0.001) (0.001) (0.000) (0.000) (0)
Panel 2: CMAP data, norm-XGR
Log-likelihood $\mathbf{-1351.9}$ $\underline{-1353.9}$ $-1362.0$ $-1354.0$ $-1438.3$ $-1439.1$ $-1469.5$ $-1440.8$ $-1463.0$ $-1439.7$ $-1426.3$
(4.697) (5.018) (3.110) (4.451) (5.834) (5.836) (6.287) (5.804) (20.05) (6.073) (0)
Accuracy 0.729 0.729 0.725 0.727 0.713 0.713 0.710 0.712 0.712 0.717 0.718
(0.003) (0.004) (0.004) (0.003) (0.003) (0.003) (0.003) (0.003) (0.002) (0.003) (0)
$F_{1}$ score 0.691 0.688 0.677 0.683 0.654 0.653 0.643 0.652 0.650 0.664 0.669
(0.005) (0.007) (0.009) (0.007) (0.006) (0.005) (0.005) (0.005) (0.002) (0.004) (0)
Strong regularity 0.888 0.857 0.706 0.815 0.998 0.998 0.929 0.997 0.426 0.999 0.998
(0.066) (0.069) (0.051) (0.068) (0.002) (0.002) (0.030) (0.004) (0.320) (0.001) (0)
Weak regularity 0.922 0.893 0.756 0.851 0.999 0.999 0.941 0.998 1.000 1.000 1.000
(0.061) (0.067) (0.051) (0.069) (0.002) (0.002) (0.026) (0.003) (0.000) (0.000) (0)
Panel 3: LTDS data, sum-XGR
Log-likelihood $\underline{-1292.0}$ $\mathbf{-1288.6}$ $\mathbf{-1288.6}$ $-1305.0$ $-1305.8$ $-1308.6$ $-1317.6$ $-1308.3$ $-1335.5$ $-1312.9$ $-1366.9$
(7.894) (9.668) (3.765) (7.316) (2.226) (2.795) (4.280) (2.670) (20.45) (2.706) (0)
Accuracy 0.729 0.729 0.727 0.724 0.732 0.730 0.728 0.730 0.725 0.736 0.730
(0.005) (0.004) (0.002) (0.004) (0.003) (0.002) (0.004) (0.002) (0.006) (0.002) (0)
$F_{1}$ score 0.727 0.728 0.725 0.721 0.728 0.726 0.723 0.726 0.721 0.733 0.726
(0.004) (0.004) (0.002) (0.006) (0.003) (0.002) (0.004) (0.002) (0.006) (0.002) (0)
Strong regularity 0.950 0.994 0.994 0.997 0.942 0.964 0.987 0.972 0.933 0.990 0.993
(0.034) (0.004) (0.009) (0.003) (0.021) (0.013) (0.008) (0.012) (0.043) (0.002) (0)
Weak regularity 0.969 0.999 0.998 1.000 0.966 1.000 1.000 1.000 1.000 1.000 1.000
(0.027) (0.001) (0.004) (0.000) (0.019) (0.000) (0.000) (0.000) (0.000) (0.000) (0)
@@
Figure 2: Individual demands as functions of driving costs (10K-Random, CMAP).